OpenStudy (studyloverhojea):

Which of the following describes how to translate the graph y = |x| to obtain the graph of y = |x - 1| - 1? 1 unit left and 1 unit down 1 unit left and 1 unit up 1 unit right and 1 unit down 1 unit right and 1 unit up

OpenStudy (lord_box):

f(x) - 1 = down f(x) + 1 = up f(x-1) = right f(x+1) = left

OpenStudy (studyloverhojea):

Okay thanks!